When one mole of nitrogen gas, N2, reacts with one mole of oxygen gas, O2, two moles of nitric

oxide, NO, are formed. In the process, 180.5 kJ of heat are required.

a. Write the thermochemical equation for the chemical reaction (Use two ways in writing the

equation)

b. What does the sign of ΔH indicate?

c. Is the reaction endothermic or exothermic? what happens to heat when the reaction takes

place?

d. If 3 moles of nitrogen gas, N2, are made to react with three moles of oxygen gas, O2, how

much heat is required in the process?

e. How much heat will be released if two moles of nitric oxide, NO, are broken down into one

mole of nitrogen gas, N2, and one mole of oxygen gas, O2?


1
Expert's answer
2022-06-07T13:18:03-0400

a.)

N2 + O2 = 2NO2 , ∆H=180,5 kJ/mol

N2 + O2 + = 2NO2 + 180,5 kJ/mol

b.)

Enthalpy change is the name given to the amount of heat evolved or absorbed in a reaction carried out at constant pressure. It is given the symbol ΔH, read as "delta H".

c.)

When enthalpy is positive and delta H is greater than zero, this means that a system absorbed heat. This is called an endothermic reaction. When enthalpy is negative and delta H is less than zero, this means that a system released heat. This is called an exothermic reaction.

N2(g) and gaseous oxygen O2(g) to form gaseous nitric oxide NO(g). This reaction is carried out at high temperature. Such reactions in which heat is absorbed by the system are called endothermic reactions.

d.)

N2 + O2 = 2NO2 , ∆H=180,5 kJ/mol

1mol N2 — 1mol O2 — 180,5 kJ

3mol N2 — 3mol O2 — X

X= 180,5 * 3 / 1 = 541,5 kJ

e.)

N2 + O2 = 2NO2 , ∆H=180,5 kJ/mol

When one mole of nitrogen gas, N2, reacts with one mole of oxygen gas, O2, two moles of nitric

oxide, NO, are formed. In the process,

180.5 kJ of heat are required.